oliviac1111:

The equation x2 – 1x – 90 = 0 has solutions {a, b}. What is a + b? –19 –9 1 10 @Florisalreadytaken

Florisalreadytaken:

\[ ax^2+bx+c \Rightarrow x^2 - 1x - 90 = 0 \] am gonna use the quadratic formula this time as well -- i dont like factoring we can see that \( a=1 \) ; \( b=-1 \) ; \( c=-90 \) the quadratic formula is in this form: \( \frac{-b\pm\sqrt{b^2-4ac}}{2a} \) can you plug the numbers n and solve for \( x_1 \) and \( x_2 \)???